15 men take 21 days of 8 hrs. each to do a piece of work. How many days of 6 hrs. each would it take for 21 women if 3 women do as much work as 2 men?

Answer by ankor@dixie-net.com(15645)

15 men take 21 days of 8 hrs. each to do a piece of work.
How many days of 6 hrs. each would it take for 21 women if 3 women do as much work as 2 men?
:
Find the no. of man-hrs to complete the job
15 * 21 * 8 = 2520 man hrs
:
2 man-hrs = 3 women-hrs
therefore
1 women-hr = 2/3 man-hr
:
let d = no. days for these women to do the job
21 * 2/3 * 6d = 2520
84d = 2520
d = 30 days
: